OpenStudy (anonymous):

One equation in a linear systemis x+2y=4. Write another equation so that the system has the unique solution (-2,3)

OpenStudy (anonymous):

y-3=1(x+3)

OpenStudy (anonymous):

sory

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

y-3=1(x+2)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

what obada is saying is that there are countless solutions to a line that passes through (-2,3)

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

all you have to do is to use y=mx+b

OpenStudy (anonymous):

slope = 1

OpenStudy (anonymous):

yes, you can set your slope to 1.. and then rock out!

OpenStudy (anonymous):

y=mx=b

OpenStudy (anonymous):

thanx!!!!!

OpenStudy (anonymous):

np. good luck!

OpenStudy (anonymous):

or you can make it perpendicular =]

OpenStudy (anonymous):

i got 2x+2 for the perp. line... is that right?